Exercise 6-20 Hunter Family Instruments makes cellos. During the past year, the company made 6,560 cellos even though the budget planned for only 5,710. The company paid its workers an average of $15 per hour, which was $1 higher than the standard labor rate. The production manager budgets four direct labor hours per cello. During the year, a total of 24,450 direct labor hours were worked. (a) Calculate the direct labor rate and efficiency variances. (If variance is zero, select "Not Applicable" and enter 0 for the amounts.) Direct labor rate variance $ Direct labor efficiency variance $
Answer #1

Direct labor rate variance = Actual Hours (Standard rate - actual rate)

= 24450 (14* - 15)

= 24450 (A)

(A) representsadverse. (Negative sign)

* = standard rate = 15 -1 = 14 (as per the language ofquestion)

Labor efficience variance = Standard rate (Standard hours for actual output - actual hours)

standard rate = 14

Actual hours = 24450

Standard hours for actual output = 6560 * 4

= 26240

Labor efficience variance = 14 ( 26240 - 24450) = 25060(F)

(A) represents negative.